OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a 12 -inch diameter pizza and an 18-inch diameter pizza. He has set the price of a -inch cheese pizza at $8.00 , based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an 18-inch cheese pizza? Show and explain your solution.

OpenStudy (anonymous):

ratio of the area goes with the square of the ratio of the diameter, because the area is a square

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so you start with \(\frac{12}{8}=\frac{3}{2}\) square to get \[(\frac{3}{2})^2=\frac{9}{4}\] and then solve \[\frac{9}{4}=\frac{x}{8}\] to find the price \(x\)

OpenStudy (anonymous):

can you help me im not really that good with fractions